(-44.6)+(49.638)+3.8-(-42.825)
1
Expert's answer
2020-03-26T12:54:50-0400

1 step: (-44.6) + (49.638) =49.638-44.6= 5.038

2 step: 5.038 + 3.8 = 8.838

3 step: 8.838 - (- 42.825) =8.838 +42.825= 51.663

Thus, (-44.6)+(49.638)+3.8-(-42.825)=51.663